# Radiative decays of the $\Omega(2012)$ as a hadronic molecule

### v1

@@ III Numerical results
We have one free parameter $\Lambda$ as introduced in the form factor of Eq. (26). Here, we take its value in the range of $900\sim 1100$ MeV as one of the $q_{\rm max}$ we have determined in Table 1. When $q_{\rm max}$ is in that range, the obtained pole position of the $\Omega(2012)$, the coupling constant $g_{\Omega^{*}\bar{K}\Xi^{*}}$, and the ratio ${\cal R}^{\bar{K}\pi\Xi}_{\bar{K}\Xi}$ are stable.
@@
With the coupling constant $\tilde{g}_{\bar{K}\Xi^{*}}=2.1$ in Table 1 and formalism in the previous section, we can obtain the radiative decay width of $\Omega(2012)\to\gamma\Omega$. The radiative decay width as a function of the parameter $\Lambda$ is shown in Fig. 4. In this work, the physical masses of the involved particles are taken from the RPP [7] and we have tabulated them in Table 2.
@@
Within the range of $\Lambda=1000\pm 100$ MeV, we obtain $\Gamma_{\Omega(2012)\to\gamma\Omega}=13.2^{+4.5}_{-3.9}$ keV. Combining this partial decay width with the theoretical total width $\Gamma_{R}=3.4$ MeV of the $\Omega(2012)$ resonance, one can estimate the branching fraction of this radiative decay process:
$BR[\Omega(2012)\to\gamma\Omega]=3.88^{+1.32}_{-1.15}\times 10^{-3}.$ (34)
Apart from the molecular picture, the radiative decay width $\Gamma_{\Omega(2012)\to\gamma\Omega}$ has also been calculated in the quark model [21], which is about $9.52$ keV. This result was derived on the basis that the $\Omega(2012)$ resonance is assigned to a $3/2^{-}$ excited 1P-wave $\Omega$ state. The partial decay width obtained here is consistent with the one obtained in Ref. [21] within uncertainties. However, it should be mentioned that the obtained branching fraction $BR[\Omega(2012)\to\gamma\Omega]$ is much larger than the one ($1.67\times 10^{-3}$) as in Ref. [21], where the total width $\Gamma_{R}=5.69$ MeV was used. In this work, we use $\Gamma_{R}=3.4$ MeV, as shown in Table 1. We hope that this decay channel will be detected in forthcoming experiments.
@@ IV Summary
The study of $\Omega$(-like) states, which contain at least three strange quarks, has long been challenging due to their scarcity. A significant advancement occurred in 2018 when the Belle Collaboration discovered the $\Omega(2012)$ resonance in the $\Xi^{0}K^{-}$ and $\Xi^{-}K^{0}_{S}$ invariant mass distributions with a combined significance of $8.3\sigma$ [1]. Subsequent analyses of its $\Xi\pi K$ and $\Xi K$ decay channels suggested its spin-parity is most likely $J^{P}=3/2^{-}$. Recent confirmations have solidified its status: the BESIII Collaboration observed it in $e^{+}e^{-}$ collisions ($3.5\sigma$) and reported a new state, the $\Omega(2109)$ ($4.1\sigma$[5], while the ALICE Collaboration observed it in $pp$ collisions with a high significance of $15\sigma$ [6]. The RPP currently lists the $\Omega(2012)$ as a three-star state with unknown spin and negative parity, reporting an average mass and width of $M_{R}=2012.5\pm 0.6$ MeV and $\Gamma_{R}=6.4^{+2.5}_{-2.0}\pm 1.6$ MeV, respectively [7].
@@
In this work, we investigate the radiative decay $\Omega(2012)\to\gamma\Omega$ within a molecular picture, where the $\Omega(2012)$ resonance with quantum numbers $J^{P}=\frac{3}{2}^{-}$, is dynamically generated by the coupled channel interactions of $\bar{K}\Xi(1530)$ and $\eta\Omega$ in $s$-wave and $\bar{K}\Xi$ in $d$-wave. By considering the dominant contribution from the triangle loop diagrams in which the $\Omega(2012)$ directly couples to $\bar{K}\Xi(1530)$ and then transitions to $\gamma\Omega$ via the exchange of a $\Xi$ hyperon, we have determined the partial decay width of $\Omega(2012)\to\gamma\Omega$ to be $13.2^{+4.5}_{-3.9}$ keV. This result offers valuable information on the nature of the $\Omega(2012)$ resonance and its decay dynamics. We anticipate that this result obtained here can be tested in future experiments by the Belle II and/or BESIII collaborations, and such experimental measurements can be used to study the nature of the $\Omega(2012)$ state in great detail.

### v2

@@ III Numerical results
We have one free parameter $\Lambda$ as introduced in the form factor of Eq. (26). Here, we take its value in the range of $850\sim 1100$ MeV as one of the $q_{\rm max}$ we have determined in Table 1. When $q_{\rm max}$ is in that range, the obtained pole position of the $\Omega(2012)$, the coupling constant $|g_{\Omega(2012)\bar{K}\Xi(1530)}|$, and the ratio ${\cal R}^{\bar{K}\pi\Xi}_{\bar{K}\Xi}$ are stable.
@@
With the coupling constant $|g_{\Omega(2012)\bar{K}\Xi(1530)}|\approx 2.1$ in Table 1 and formalism in the previous section, we can obtain the radiative decay width of $\Omega(2012)\to\gamma\Omega$. The radiative decay width as a function of the parameter $\Lambda$ is shown in Fig. 4. In this work, the physical masses of the involved particles are taken from the RPP [37] and we have tabulated them in Table 2.
@@
Within the range of $\Lambda=1000^{+100}_{-150}$ MeV, we obtain $\Gamma_{\Omega(2012)\to\gamma\Omega}=13.2^{+4.5}_{-5.6}$ keV. Combining this partial decay width with the theoretical total width $\Gamma_{R}=3.4\pm 0.3$ MeV of the $\Omega(2012)$ resonance, one can estimate the branching fraction of this radiative decay process:
$BR[\Omega(2012)\to\gamma\Omega]=3.9^{+1.4}_{-1.7}\times 10^{-3}.$ (34)
Apart from the molecular picture, the radiative decay width $\Gamma_{\Omega(2012)\to\gamma\Omega}$ has also been calculated in the quark model [29], which is about $9.52$ keV. This result was derived on the basis that the $\Omega(2012)$ resonance is assigned to a $3/2^{-}$ excited 1P-wave $\Omega$ excited state. The partial decay width obtained here is consistent with the one obtained in Ref. [29] within uncertainties. However, it should be noted that the branching fraction obtained here, $BR[\Omega(2012)\to\gamma\Omega]$, differs from the value ($1.67\times 10^{-3}$) given in Ref. [29], where a total width $\Gamma_{R}=5.69$ MeV was used. In this work, we employ $\Gamma_{R}=3.4\pm 0.3$ MeV, as listed in Table I. Uncertainties arising from the errors in its coupling constant to the $\bar{K}\Xi(1530)$ channel are not included, since their contributions are small compared to the error originating from the range of the cutoff parameter $\Lambda$. We hope that this decay channel will be measured in future experiments, providing a test for the model calculations presented here.
@@ IV Summary
The study of $\Omega$(-like) states, which contain at least three strange quarks, has long been challenging due to their scarcity. A significant advancement occurred in 2018 when the Belle Collaboration discovered the $\Omega(2012)$ resonance in the $\Xi^{0}K^{-}$ and $\Xi^{-}K^{0}_{S}$ invariant mass distributions with a combined significance of $8.3\sigma$ [64]. Subsequent analyses of its $\Xi\pi\bar{K}$ and $\Xi\bar{K}$ decay channels suggested its spin-parity is most likely $J^{P}=3/2^{-}$. Recent confirmations have solidified its status: the BESIII Collaboration observed it in $e^{+}e^{-}$ collisions ($3.5\sigma$) and reported a new state, the $\Omega(2109)$ ($4.1\sigma$[1], while the ALICE Collaboration observed it in $pp$ collisions with a high significance of $15\sigma$ [2]. The RPP currently lists the $\Omega(2012)$ as a three-star state with unknown spin and negative parity, reporting an average mass and width of $M_{R}=2012.5\pm 0.6$ MeV and $\Gamma_{R}=6.4^{+2.5}_{-2.0}\pm 1.6$ MeV, respectively [37].
@@
In this work, we investigate the radiative decay $\Omega(2012)\to\gamma\Omega$ within a molecular picture, where the $\Omega(2012)$ resonance with quantum numbers $J^{P}=\frac{3}{2}^{-}$, is dynamically generated by the coupled channel interactions of $\bar{K}\Xi(1530)$ and $\eta\Omega$ in $s$-wave and $\bar{K}\Xi$ in $d$-wave. By considering the dominant contribution from the triangle loop diagrams in which the $\Omega(2012)$ directly couples to $\bar{K}\Xi(1530)$ and then transitions to $\gamma\Omega$ via the exchange of a $\Xi$ hyperon, we have determined the partial decay width of $\Omega(2012)\to\gamma\Omega$ to be $13.2^{+4.5}_{-5.6}$ keV. This result offers valuable information on the nature of the $\Omega(2012)$ resonance and its decay dynamics. We anticipate that this result obtained here can be tested in future experiments by the Belle II and/or BESIII collaborations, and such experimental measurements can be used to study the nature of the $\Omega(2012)$ state in great detail.
@@ Acknowledgements.
DATA AVAILABILITY
The data support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.